OpenStudy (anonymous):

A line has slope -4 and y-intercept (0, -2). What is the slope-intercept form of the equation of this line? A. y = -2x + 4 B. y = -2x - 4 C. y = -4x - 2 D. y = -4x + 2 Someone please help!

OpenStudy (daniellelovee):

y=mx+b m=slope b=yint

OpenStudy (daniellelovee):

there is only one answer that has the correct y-intercept do you know which one is it now?

OpenStudy (anonymous):

Well you can eliminate A and B

OpenStudy (anonymous):

Is it D???

OpenStudy (daniellelovee):

close but no because the y intercept is -2 not +2

OpenStudy (anonymous):

But if its a negative isn't it suppose to become a positive?

OpenStudy (anonymous):

even if its the y intercept.

OpenStudy (daniellelovee):

oh i see where you get confused because the formula says y=mx+b no the +b can be -b as well it depends on where the intercept is and is this case is negative

OpenStudy (anonymous):

So then it would be C..?

OpenStudy (daniellelovee):

yup

OpenStudy (anonymous):

It's C cuz the slope (m) =-4 and the y-intercept (b)=-2 the slope intercept form is y=mx+b apply the givens y=-4x-2

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (daniellelovee):

no problem :)